OpenStudy (anonymous):

A(t)=(t/t+1)-1 A(1/11) = ?

OpenStudy (anonymous):

A(t) = (t/(t+1))-1 A(1/11) = ((1/11)/(12/11))-11/11 =-11/12

OpenStudy (anonymous):

i'm confused on how we got -11/12

OpenStudy (anonymous):

why was it switched? and how did we get there?

OpenStudy (anonymous):

Just substitute t = 1/11.

OpenStudy (anonymous):

i did but my math does not add up -_-

OpenStudy (anonymous):

I don't get the -11/12

OpenStudy (anonymous):

I don't understand why we flipped the fraction